Compute and interpret the correlation coefficient for the aptitude scores and grade point averages in Exercise 8 on page 373.



1
Expert's answer
2021-09-20T16:20:27-0400

Since, the data is incomplete, we assume as follows:

Compute and interpret the correlation coefficient for the aptitude scores and grade point averages given as:

Aptitude scores (X) (out of 100) : 78, 85, 92, 98, 52

Grade point average (Y) (out of 10): 8, 9, 9, 9, 5

Solution:


X Values

∑ = 405

Mean = 81

∑(X - Mx)2 = SSx = 1276


Y Values

∑ = 40

Mean = 8

∑(Y - My)2 = SSy = 12


X and Y Combined

N = 5

∑(X - Mx)(Y - My) = 119


R Calculation

r = ∑((X - My)(Y - Mx)) / "\sqrt{(SS_x)(SS_y)}"

r = 119 / "\sqrt{(1276)(12)}" = 0.9617


This is a strong positive correlation, which means that high X variable scores go with high Y variable scores (and vice versa).